OpenStudy (anonymous):

Integration By Parts Hello, I'm actually having problems with the algebra on this Calculus question. The problem is; \[\int\limits_{}^{}e^{7x}*\cos(9x)dx\] So I get all of the Calculus right (confirmed) and get to; \[\int\limits_{}^{}(1/7)\cos(9x)e^{7x} + (9/7)\sin(9x)e^{7x} - (81/49)\int\limits_{}^{}e^{7x}\cos(9x)\] So we have the original equation on both sides and need to add that to both sides. This is where I seem to have problems. I know the answer has a denominator of 130, but I keep getting 81. Any help would be greatly appreciated.

OpenStudy (amistre64):

the last is the same "kinda" as the first

OpenStudy (amistre64):

A = b-2A ; solve for b

OpenStudy (amistre64):

i should finish reading before commentling lol

OpenStudy (anonymous):

e^(7x) cos(9x) = I Rest = X X - (81/49)I = I X = I + (81/49)I X / (81/49) = I + I X / (81/49) = 2I X / (81/49) / 2 = I

OpenStudy (amistre64):

cos(9x) + e^7x sin(9x)/9 - 7 e^7x -cos(9x)/81 + 49 e^7x ............ {S} e^7x cos(9x)=e^7x sin(9x)/9 +7 e^7x cos(9x)/81 -{S} 49 e^7x cos(9x)/81dx {S} e^7x cos(9x)=e^7x sin(9x)/9 +7 e^7x cos(9x)/81 -49/81{S} e^7x cos(9x)dx +49/81{S} e^7x cos(9x)dx +49/81{S} e^7x cos(9x)dx ------------------------------------------------------------------- 130{S} e^7x cos(9x) = e^7x sin(9x)/9 +7 e^7x cos(9x)/81 \130 \130 maybe

OpenStudy (anonymous):

But how did that 49/81 turn into 130? It feels like magic!

OpenStudy (amistre64):

so if i dealt with my signs correctly, and i never do .... e^7x (sin(9x/9) + 7 cos(9x)/81) --------------------------- 130

OpenStudy (amistre64):

A = b - 49/81 A add 49/81 A to both sides

OpenStudy (amistre64):

if my signs are right lol

OpenStudy (anonymous):

Right, so then you've got A + (49/81)A = B Divide by 49/81 on both sides to eliminate, and you've got A + A = B/(49/81)

OpenStudy (anonymous):

Or am I just out to lunch with my algebra here? I did the same A = b - 49/81A up above with X and I.

OpenStudy (amistre64):

your algebra is off since you cant divide out of a sum

OpenStudy (amistre64):

i see what i did; 130/81; then multiply thru by 81/130 to clear the left side

OpenStudy (amistre64):

all I did was divide off the 130 :)

OpenStudy (anonymous):

But where did 130/81 come from?

OpenStudy (anonymous):

49 turned into 130 by adding 81 to it? Where'd the 81 come to add to it?

OpenStudy (amistre64):

from the mess that you get at the end 1A = stuff - 49/81 A ; ADD 49/81 A to both sides 1A + 49/81 A = stuff factor out the A A(1+ 49/81) = stuff since 1 + 49/81 = 130/81, multiply thru by 81/130 A = 81/130 * stuff

OpenStudy (amistre64):

the 81 is the denominator of the constant at the end of all that mess

OpenStudy (amistre64):

49/81 is the pulled out constant from your integrations

OpenStudy (anonymous):

Oh man.. such a silly freaking mistake.

OpenStudy (amistre64):

{S} e^7x cos(9x) dx = e^7x sin(9x)/9 - {S} 7 e^7x sin(9x)/9 dx ... whew!! glad you see it lol

OpenStudy (anonymous):

All from missing the factoring. Amazing how it can just destroy your answer.

OpenStudy (amistre64):

yeah, its fun on a test :)

OpenStudy (anonymous):

Haha, thats great man. Thanks so much for the help! You've helped me a ton on here :) I hope you're becoming a math professor!! Off to work more homework!

OpenStudy (amistre64):

good luck :)